Burke Tires just paid a dividend of D0 = $1.30. Analysts expect the company's dividend to grow by 24% this year, by 17% in Year 2, grow by 10% in Year 3 and at a constant rate of 6% in Year 4 and thereafter. The required return on this low-risk stock is 10.00%. What is the best estimate of the stock's current market value?
Answer #1

D0 = $1.30
D1 = $1.30*1.24 = $1.612
D2 = $1.612*1.17 = $1.886
D3 = $1.886*1.10 = $2.075
D4 = $2.075*1.06 = $2.199

Constant Growth rate, g = 6%
Required return , r = 10%

P3 = D4 / (r - g) = 2.199 / (0.10 - 0.06) = $54.978

P0 = 1.612 /1.10 + 1.886 / 1.10^2 + 2.075 / 1.10^3 + 54.978 / 1.10^3 = $45.89

So, the current value (intrinsic) of the stock is $45.89